A turntable that spins at a constant 75.0 rpmtakes 4.00 s to reach this angular speed after it is turned on. Find its angular acceleration (in rad/s2), assuming it to be constant, and the number of degrees it turns through while speeding up.
Answer #1

angular acceleration(a)=(75*2*pi/60 -0 )/4=5*pi/8 =1.963rad/s2

number of degrees turned= s;

v2=u2+2*a*s

(75*2*pi/60)2= 0+2*5*pi/8 * s

angle rotated=5*pi= 180*5=9000